OpenStudy (j.f):

Metal g^-1(x) Show the steps it takes to arrive at the answer (x+1)^1/2 If I'm late to respond its because of my computer

OpenStudy (freckles):

I don't get it. How can we arrive at (x+1)^(1/2) if we aren't given anything? Is metal a math terminology? If so what does it mean to metal g^(-1)(x)? Or maybe g(x)=(x+1)^(1/2) and you want to find the g^(-1)(x)?

OpenStudy (freckles):

\[y=(x+1)^\frac{1}{2}\] square both sides first and solve for x

OpenStudy (j.f):

AHhhh, brain fart! f(x)=2x+1 and g(x)=x^2-1

OpenStudy (freckles):

we only need g(x)=x^2-1 to find g^(-1)(x)

OpenStudy (freckles):

y=x^2-1 add 1 on both sides as a first step

OpenStudy (freckles):

you are trying to solve for x

OpenStudy (freckles):

please show what you have after doing that step

OpenStudy (j.f):

y+(1)=x^2-1+(1) is this properly set up?

OpenStudy (freckles):

y+1=x^2 correct

OpenStudy (freckles):

now you can take the square root of both sides ! if you assume x>0 you will have your answer

OpenStudy (j.f):

wouldn't the problem look like y+1=x^2= y+1/x=x? where is the 1/2 coming into play? Sorry for my ignorance I just haven't done these in some time

OpenStudy (freckles):

I don't know what you did to get all of that

OpenStudy (freckles):

\[y+1=x^2\] I know to solve for x you take square root of both sides and if you assume x>0 you have your answer

OpenStudy (freckles):

you do know \[\sqrt{9}=\sqrt{3^2}=3\] you see that the square root and the square cancel right?

OpenStudy (j.f):

Yes

OpenStudy (freckles):

Example below: just like if you are solving \[x^2=9 \\ \text{ you would take square root of both sides } \\ x=\sqrt{9}=3 \text{ or } x=-\sqrt{9}=-3\]

OpenStudy (freckles):

here you do the same thing \[x^2=y+1\] \[\text{ if } x>0 \text{ then } x=\sqrt{y+1} \\ \\ \text{ if } x<0 \text{ then } x=-\sqrt{y+1}\]

OpenStudy (freckles):

now you interchange the variables and after doing that you can call y g^(-1)(x) if you choose

OpenStudy (j.f):

I see. (1+x)^1/2 is the same as sqrt x+1

OpenStudy (freckles):

yes

OpenStudy (j.f):

danke